# The Ginsburg–Sands theorem and computability theory

### v1

@@ 3.1. Spaces and subspaces
    for every $n\in\mathbb{N}$, $U_{n}=\bigcap_{n\in F}V_{n}$, where $F$ is the finite set coded by $n$.
@@ 3.3. Indiscrete and discrete spaces
The implication from (1) to (2) is straightforward, and (2) implies (3) trivially. To prove (3) implies (1), we argue in $\mathsf{RCA}_{0}$ and assume (3). To begin, we derive $\mathsf{B}\Sigma^{0}_{2}$, which we take in the guise of $(\forall m)\mathsf{RT}^{1}_{m}$ using Hirst’s theorem. So fix $c:\mathbb{N}\to m$ for some $m\in\mathbb{N}$, and suppose towards a contradiction that $c$ has no infinite homogeneous set. For all $x,s\in\mathbb{N}$, define $V_{\langle x,s\rangle}=\{x\}\cup\{z>s:c(z)=c(x)\}$, and let $\langle\mathbb{N},\mathcal{U},k\rangle$ be the CSC space generated by $\langle V_{n}:n\in\mathbb{N}\rangle$. By assumption, $\{z:c(z)=c(x)\}$ is bounded by some $s$, meaning that $V_{\langle x,s\rangle}=\{x\}$. Thus, $\langle\mathbb{N},\mathcal{U},k\rangle$ is an infinite discrete space. Now let $Y\subseteq\mathbb{N}$ be an infinite effectively discrete subspace of $\langle\mathbb{N},\mathcal{U},k\rangle$, with witness $d:Y\to\mathbb{N}$. That is, if we write $\mathcal{U}=\langle U_{n}:n\in\mathbb{N}\rangle$ then $U_{d(x)}\cap Y=\{x\}$ for all $x\in Y$. We define a sequence $x_{0}<\cdots<x_{m}$ in $Y$ as follows. Let $x_{0}\in Y$ be arbitrary, and suppose we have defined $x_{i}$ for some $i<m$. By Proposition 3.3 (2), we can decode $d(x_{i})$ as a nonempty finite set $F$ such that $U_{d(x_{i})}=\bigcap_{n\in F}V_{n}$. Now, each $n\in F$ is the form $\langle z,s\rangle$ for some $z$ and $s$, and since $x_{i}\in V_{\langle z,s\rangle}$ it follows that every $y>s$ with $c(y)=c(x_{i})$ also belongs to $V_{\langle z,s\rangle}$. Choose the largest $s$ such that $\langle z,s\rangle\in F$ for some $z$ and let $x_{i+1}$ be the least element of $Y$ larger than $x_{i}$ and $s$. This completes the definition. By induction on $j\leq m$, $c(x_{i})\neq c(y)$ for all $i<j$ and all $y\geq x_{j}$, so in particular, $c(x_{i})\neq c(x_{j})$. Hence, the map $x_{i}\mapsto c(x_{i})$ is an injection of a finite set of size $m+1$ into $m$, which contradicts $\mathsf{I}\Sigma^{0}_{1}$. We conclude that $c$ has an infinite homogeneous set after all, and hence that $\mathsf{B}\Sigma^{0}_{2}$ holds. We now use (3) and $\mathsf{B}\Sigma^{0}_{2}$ to derive $\mathsf{ACA}_{0}$. To this end, it suffices to prove that the range of every one-to-one function $f:\mathbb{N}\to\mathbb{N}$ exists. Fix such an $f$. For $x,s\in\mathbb{N}$, let $\operatorname{ran}(f)\operatorname{\upharpoonright}x~{}[s]=\operatorname{ran}(f\operatorname{\upharpoonright}s+1)\operatorname{\upharpoonright}x=\{w<x:(\exists y\leq s)[f(y)=w]\}$, and define
@@ 3.4. The initial segment and final segment topologies
To see that (1) implies (2), let $\langle X,\mathcal{U},k\rangle$ be a CSC space with the weak initial segment topology. We define $h:\mathbb{N}\to X$ as follows. By properties (1b) and (1c), $\mathsf{ACA}_{0}$ can find a sequence of sets $U_{0}\subset U_{1}\subset\cdots$ with $|U_{i}|=i+1$ for all $i\in\mathbb{N}$. By property (1b), each finite $U\in\mathcal{U}$ is equal to $U_{i}$ for some $i$, so by property (1d), $X=\bigcup_{i\in\mathbb{N}}U_{i}$. Define $h:\mathbb{N}\to X$ by letting $h(i)$ be the unique element of $U_{i}\smallsetminus U_{j<i}U_{j}$. Thus, $h$ is bijective, and by induction, $U_{j}=\{h(i):i\leq j\}$ for each $j$. Now fix any $U\in\mathcal{U}$ other than $\emptyset$ and $X$. By property (1a), $U$ is finite, hence equal to $U_{j}$ for some $j$, which means $U=\{h(i):i\leq j\}$. Thus, $h$ is a homeomorphism witnessing that $X$ has the initial segment topology. A similar argument shows that (1) implies (3). We now show that (2) implies (1); the proof that (3) implies (1) is similar. So, assume (2), and fix a one-to-one $f:\mathbb{N}\to\mathbb{N}$ with the aim of showing that $\operatorname{ran}(f)$ exists. By bounded $\Sigma^{0}_{1}$ comprehension, $\operatorname{ran}(f)$ is unbounded. For each $x$, define
@@
A straightforward induction in $\mathsf{RCA}_{0}$ shows that $V_{n}\subseteq V_{n+1}$ and $|V_{n+1}\smallsetminus V_{n}|\leq 1$ for all $n$. Since $|V_{0}|=1$, it follows that all the $V_{n}$ are finite. We claim that for each $s>0$ there is an $n$ such that $|V_{n}|=s$. Suppose not, and fix an $s>0$ witnessing this. Since $\operatorname{ran}(f)$ is unbounded, there exist $x_{0}<\cdots<x_{s}\in\operatorname{ran}(f)$, and because $f$ is injective this means $|V_{2x_{s}+1}|\geq|\{y:(\exists i\leq s)[f(y)=x_{i}]\}|>s$. Thus, $\mathsf{RCA}_{0}$ can prove that there is a least $n$ such that $|V_{n}|>s$. Obviously we must have $s>1$ and so $n\geq 1$, and by assumption, $|V_{n-1}|<s$. But then $|V_{n}|>|V_{n-1}|+1$, a contradiction. This proves the claim. Now let $\langle\mathbb{N},\mathcal{U},k\rangle$ be the CSC space generated by $\langle V_{n}:n\in\mathbb{N}\rangle$. The preceding argument implies that this space has the weak initial segment topology. So apply (2) to get a homeomorphism $h:\mathbb{N}\to\mathbb{N}$, and fix any $x$. Since $V_{2x}$ and $V_{2x+1}$ are finite, we can find the least $j_{0}$ and $j_{1}$ such that $h(j_{0}+1)\notin V_{2x}$ and $h(j_{1}+1)\notin V_{2x+1}$, which means that $V_{2x}=\{h(i):i\leq j_{0}\}$ and $V_{2x+1}=\{h(i):i\leq j_{1}\}$. By construction, $j_{0}\leq j_{1}\leq j_{0}+1$, and we have that $x\in\operatorname{ran}(f)$ if and only if $j_{1}=j_{0}+1$. ∎
@@
We prove (1); the proof of (2) is analogous. Fix $\langle X,\mathcal{U},k\rangle$ satisfying (1a)–(1d) in Definition 3.12. (In fact, we will not need to use (1d).) We define sequences $x_{0},x_{1},\ldots\in X$ and $U_{0},U_{1},\ldots\in\mathcal{U}$ such that $x_{i}\in U_{j}$ if and only if $i\leq j$. We claim that $Y=\{x_{0},x_{1},\ldots\}$ is a subspace of $X$ with the initial segment topology, as witnessed by the map $h:i\mapsto x_{i}$. Note that $U_{j}\cap Y=\{x_{i}:i\leq j\}$ for each $j$, so we wish to show that the basic open sets in $Y$ in addition to $\emptyset$ and $Y$ are precisely the sets $U_{j}\cap Y$ for some $j$. Since $U_{j}\in\mathcal{U}$, it is certainly true that $U_{j}\cap Y$ is basic open in $Y$. Conversely, consider any basic open set in $Y$ different from $\emptyset$ and $Y$. Then this equals $U\cap Y$ for some $U\in\mathcal{U}$ different from $\emptyset$ and $X$. By property (1a), $U$ is finite, so we can fix the largest $j$ such that $x_{i}\in U$ for all $i\leq j$. In particular, $U_{j}\cap Y\subseteq U\cap Y$. Since $x_{j+1}\in U_{j+1}\smallsetminus U$, we cannot have $U_{j+1}\subseteq U$, and so by property (1b) it must be that $U\subseteq U_{j+1}$. Moreover, since $x_{j+1}\notin U$ we in fact have that $U\subseteq U_{j+1}\smallsetminus\{x_{j+1}\}$. Thus, $U\cap Y\subseteq(U_{j+1}\smallsetminus\{x_{j+1}\})\cap Y=U_{j}\cap Y$, as desired. It remains to show how to construct the sequences $x_{0},x_{1},\ldots$ and $U_{0},U_{1},\ldots$, which we do by primitive recursion. Fix $j$, and suppose we have defined $x_{i}$ and $U_{i}$ for all $i<j$. We claim there is a finite $U\in\mathcal{U}$ and an $x\in U$ such that $x_{i}\in U$ and $x\notin U_{i}$ for all $i<j$. Indeed, using bounded $\Pi^{0}_{1}$ comprehension we can form the set $\{s_{i}:i<j\}$ such that $|U_{i}|=s_{i}$ for all $i<j$. Let $s=\sum_{i<j}s_{i}$, and using property (1c) let $U\in\mathcal{U}$ be a finite set with at least $s+1$ many elements. Since $\mathsf{I}\Sigma^{0}_{1}$ proves that there is no injection from a larger finite set into a strictly smaller one, we cannot have $U\subseteq U_{i}$ for any $i<j$. Hence, $\bigcup_{i<j}U_{i}\subseteq U$ by property (1b), and this containment must be strict. Fixing any $x\in U\smallsetminus\bigcup_{i<j}U_{i}$ completes the proof of the claim. Now by property (1a), being a finite element of $\mathcal{U}$ is definable by a $\Sigma^{0}_{1}$ formula (with parameters). Thus, the least $U$ and $x\in U$ as in the claim are definably by a $\Sigma^{0}_{1}$ formula. We let $U_{j+1}=U$ and $x_{j+1}=x$. This completes the recursive definition and completes the proof. ∎
@@
Write $\mathcal{U}=\langle U_{n}:n\in\mathbb{N}\rangle$. Suppose $X$ has the initial segment topology; the case where it has the final segment topology is symmetric. Let $h:\mathbb{N}\to X$ be the witnessing homeomorphism, and consider any open code $O$. If $O_{\mathcal{U}}=X$ then it clearly belongs to $\mathcal{U}$, so suppose otherwise. Then every $U_{n}$ for $n\in\operatorname{ran}(O)$ is either $\emptyset$ or has the form $\{h(i):i\leq j_{n}\}$ for some $j_{n}$, and the numbers $j_{n}$ must be bounded else $O_{\mathcal{U}}$ would again be $X$. So fix $m$ such that $j_{n}\leq j_{m}$ for all $n\in\operatorname{ran}(O)$ for which $j_{n}$ is defined. Then $O_{\mathcal{U}}=U_{m}\in\mathcal{U}$. ∎
@@ 4. General spaces
For the implication from (1) to (2), we note that the proof of the Ginsburg–Sands theorem given in Section 2 readily formalizes in $\mathsf{ACA}_{0}$ when applied to CSC spaces. The implication from (2) to (3) is trivial. It thus remains to show that (3) implies (1). To this end, we argue in $\mathsf{RCA}_{0}$ and assume $\mathsf{wGS}$. Note that $\mathsf{wGS}$ trivially implies $\mathsf{wGS}^{\operatorname{cl}}$, which we in turn prove is equivalent to $\mathsf{GST}_{1}$ in Theorem 4.8 below. Moreover, in Corollary 6.10 we show that $\mathsf{GST}_{1}$ implies $\mathsf{RT}^{2}_{2}$, and hence $\mathsf{B}\Sigma^{0}_{2}$. Therefore, we may avail ourselves of $\mathsf{B}\Sigma^{0}_{2}$ in our argument. We prove that the range of a given one-to-one function $f:\mathbb{N}\to\mathbb{N}$ exists. Given $x,s\in\mathbb{N}$, using the notation from the proof of Proposition 3.10, define
@@
and let $\langle\mathbb{N},\mathcal{U},k\rangle$ be the CSC space generated by $\langle V_{n}:n\in\mathbb{N}\rangle$. We claim that if a number $x$ is sufficiently large with respect to a number $w$ then $w$ belongs to every basic open set containing $x$. To see this, fix $w$ and choose $x>w$ so that $\operatorname{ran}(f)\operatorname{\upharpoonright}w=\operatorname{ran}(f)\operatorname{\upharpoonright}w~{}[x]$, which exists by $\mathsf{B}\Sigma^{0}_{2}$. Fix $U\in\mathcal{U}$ containing $x$. Then $U$ is a finite intersection of sets of the form $V_{\langle z,s\rangle}$ for some $z$ and $s$, and since no element of $V_{\langle z,s\rangle}$ can be larger than $z$, we necessarily have $z\geq x$. Consider such a $V_{\langle z,s\rangle}$. If $x=z$, then $w\in V_{\langle z,s\rangle}=V_{\langle x,s\rangle}$ since $\operatorname{ran}(f)\operatorname{\upharpoonright}w~{}[t]=\operatorname{ran}(f)\operatorname{\upharpoonright}w$ for all $t\geq x$ by choice of $x$. If $x<z$ then we must have $\operatorname{ran}(f)\operatorname{\upharpoonright}x~{}[z]=\operatorname{ran}(f)\operatorname{\upharpoonright}x~{}[z+s]$, and hence also $\operatorname{ran}(f)\operatorname{\upharpoonright}w~{}[z]=\operatorname{ran}(f)\operatorname{\upharpoonright}w~{}[z+s]$ since $w<x$, so $w\in V_{\langle z,s\rangle}$. This proves the claim. Apply $\mathsf{wGS}$ to $X$ to obtain an infinite subspace $Y$. By the preceding claim, we immediately know that $Y$ cannot be $T_{1}$ nor have the final segment topology. And since, for $w<x$, we have that $V_{\langle w,0\rangle}$ contains $w$ but not $x$, we know that $Y$ cannot be indiscrete. This means that $Y$ has the initial segment topology. There is thus a bijection $h:\mathbb{N}\to Y$ such that the basic open sets of $Y$ in the subspace topology are precisely $\{h(i):i\leq j\}$ for some $j\in\mathbb{N}$. Let $Y=\{y_{0}<y_{1}<\cdots\}$; we claim that $h(n)=y_{n}$ for all $n$. To see this, first suppose that $h(n)>h(m)$ for some $n<m$, and fix $j$ such that $V_{\langle h(m),0\rangle}\cap Y=\{h(i):i\leq j\}$. Then $V_{\langle h(m),0\rangle}\cap Y$ contains $h(n)$, contradicting the fact that this intersection only contains $h(m)$ and otherwise numbers smaller than $y_{m}$. It follows that $h(n)<h(m)$ for all $n<m$, whence an easy induction using the fact that $h$ is bijective shows that $h(n)=y_{n}$. This proves the claim. In particular, every basic open set in $X$ is closed downward in $Y$ under $\leq$. We now use $Y$ to prove the existence of $\operatorname{ran}(f)$. Fix $w\in\mathbb{N}$. Since $Y$ is infinite, we can find $z>x>w$ with $x,z\in Y$. We must then have that $\operatorname{ran}(f)\operatorname{\upharpoonright}x=\operatorname{ran}(f)\operatorname{\upharpoonright}x~{}[z]$, and hence also that $\operatorname{ran}(f)\operatorname{\upharpoonright}w=\operatorname{ran}(f)\operatorname{\upharpoonright}w~{}[z]$. If not, then there would be an $s$ such that $\operatorname{ran}(f)\operatorname{\upharpoonright}x~{}[z]\neq\operatorname{ran}(f)\operatorname{\upharpoonright}x~{}[z+s]$. But then $V_{\langle z,s\rangle}$ would be a basic open set in $X$ containing $z$ but not $x<z$, even though both belong to $Y$. This completes the proof. ∎

### v2

@@ 3.1. Spaces and subspaces
    for every $m\in\mathbb{N}$, $U_{m}=\bigcap_{n\in F}V_{n}$, where $F$ is the finite set coded by $m$.
@@ 3.3. Indiscrete and discrete spaces
The implication from (1) to (2) is straightforward, and (2) implies (3) trivially. To prove (3) implies (1), we argue in $\mathsf{RCA}_{0}$ and assume (3). To begin, we derive $\mathsf{B}\Sigma^{0}_{2}$, which we take in the guise of $(\forall m)\mathsf{RT}^{1}_{m}$ using Hirst’s theorem. So fix $c:\mathbb{N}\to m$ for some $m\in\mathbb{N}$, and suppose towards a contradiction that $c$ has no infinite homogeneous set. For all $x,s\in\mathbb{N}$, define $V_{\langle x,s\rangle}=\{x\}\cup\{z>s:c(z)=c(x)\}$, and let $\langle\mathbb{N},\mathcal{U},k\rangle$ be the CSC space generated by $\langle V_{n}:n\in\mathbb{N}\rangle$. By assumption, $\{z:c(z)=c(x)\}$ is bounded by some $s$, meaning that $V_{\langle x,s\rangle}=\{x\}$. Thus, $\langle\mathbb{N},\mathcal{U},k\rangle$ is an infinite discrete space. Now let $Y\subseteq\mathbb{N}$ be an infinite effectively discrete subspace of $\langle\mathbb{N},\mathcal{U},k\rangle$, with witness $d:Y\to\mathbb{N}$. That is, if we write $\mathcal{U}=\langle U_{n}:n\in\mathbb{N}\rangle$ then $U_{d(x)}\cap Y=\{x\}$ for all $x\in Y$. We define a sequence $x_{0}<\cdots<x_{m}$ in $Y$ as follows. Let $x_{0}\in Y$ be arbitrary, and suppose we have defined $x_{i}$ for some $i<m$. By Proposition 3.3 (2), we can decode $d(x_{i})$ as a nonempty finite set $F$ such that $U_{d(x_{i})}=\bigcap_{n\in F}V_{n}$. Now, each $n\in F$ is the form $\langle z,s\rangle$ for some $z$ and $s$, and since $x_{i}\in V_{\langle z,s\rangle}$ it follows that every $y>s$ with $c(y)=c(x_{i})$ also belongs to $V_{\langle z,s\rangle}$. Choose the largest $s$ such that $\langle z,s\rangle\in F$ for some $z$ and let $x_{i+1}$ be the least element of $Y$ larger than $x_{i}$ and $s$. This completes the definition. By induction on $j\leq m$, $c(x_{i})\neq c(y)$ for all $i<j$ and all $y\geq x_{j}$, so in particular, $c(x_{i})\neq c(x_{j})$. Hence, the map $x_{i}\mapsto c(x_{i})$ is an injection of a finite set of size $m+1$ into $m$, which contradicts $\mathsf{I}\Sigma^{0}_{1}$. We conclude that $c$ has an infinite homogeneous set after all, and hence that $\mathsf{B}\Sigma^{0}_{2}$ holds.
We now use (3) and $\mathsf{B}\Sigma^{0}_{2}$ to derive $\mathsf{ACA}_{0}$. To this end, it suffices to prove that the range of every one-to-one function $f:\mathbb{N}\to\mathbb{N}$ exists. Fix such an $f$. For $x,s\in\mathbb{N}$, let $\operatorname{ran}(f)\operatorname{\upharpoonright}x~{}[s]=\operatorname{ran}(f\operatorname{\upharpoonright}s+1)\operatorname{\upharpoonright}x=\{w<x:(\exists y\leq s)[f(y)=w]\}$, and define
@@ 3.4. The initial segment and final segment topologies
To see that (1) implies (2), let $\langle X,\mathcal{U},k\rangle$ be a CSC space with the weak initial segment topology. We define $h:\mathbb{N}\to X$ as follows. By properties (1b) and (1c), $\mathsf{ACA}_{0}$ can find a sequence of sets $U_{0}\subset U_{1}\subset\cdots$ with $|U_{i}|=i+1$ for all $i\in\mathbb{N}$. By property (1b), each finite $U\in\mathcal{U}$ is equal to $U_{i}$ for some $i$, so by property (1d), $X=\bigcup_{i\in\mathbb{N}}U_{i}$. Define $h:\mathbb{N}\to X$ by letting $h(i)$ be the unique element of $U_{i}\smallsetminus U_{j<i}U_{j}$. Thus, $h$ is bijective, and by induction, $U_{j}=\{h(i):i\leq j\}$ for each $j$. Now fix any $U\in\mathcal{U}$ other than $\emptyset$ and $X$. By property (1a), $U$ is finite, hence equal to $U_{j}$ for some $j$, which means $U=\{h(i):i\leq j\}$. Thus, $h$ is a homeomorphism witnessing that $X$ has the initial segment topology. A similar argument shows that (1) implies (3).
We now show that (2) implies (1); the proof that (3) implies (1) is similar. So, assume (2), and fix a one-to-one $f:\mathbb{N}\to\mathbb{N}$ with the aim of showing that $\operatorname{ran}(f)$ exists. By bounded $\Sigma^{0}_{1}$ comprehension, $\operatorname{ran}(f)$ is unbounded. For each $x$, define
@@
A straightforward induction in $\mathsf{RCA}_{0}$ shows that $V_{n}\subseteq V_{n+1}$ and $|V_{n+1}\smallsetminus V_{n}|\leq 1$ for all $n$. Since $|V_{0}|=1$, it follows that $|V_{n}|\leq n+1$ for all $n$, and so in particular that each $V_{n}$ is finite. We claim that for each $s>0$ there is an $n$ such that $|V_{n}|=s$. Indeed, since $\operatorname{ran}(f)$ is unbounded there exist $x_{0}<\cdots<x_{s}\in\operatorname{ran}(f)$. Because $f$ is injective, this means $|V_{2x_{s}+1}|\geq|\{y:(\exists i\leq s)[f(y)=x_{i}]\}|>s$. Thus, $\mathsf{RCA}_{0}$ can prove that there is a least $n$ such that $|V_{n}|>s$. Obviously we must have $s>1$ and so $n\geq 1$. By minimality of $n$, we have $|V_{n-1}|\leq s$, hence by our observation above $|V_{n-1}|=|V_{n}|-1$. It follows that $|V_{n-1}|=s$, which proves the claim. Now let $\langle\mathbb{N},\mathcal{U},k\rangle$ be the CSC space generated by $\langle V_{n}:n\in\mathbb{N}\rangle$. The preceding argument implies that this space has the weak initial segment topology. So apply (2) to get a homeomorphism $h:\mathbb{N}\to\mathbb{N}$, and fix any $x$. Since $V_{2x}$ and $V_{2x+1}$ are finite, we can find the least $j_{0}$ and $j_{1}$ such that $h(j_{0}+1)\notin V_{2x}$ and $h(j_{1}+1)\notin V_{2x+1}$, which means that $V_{2x}=\{h(i):i\leq j_{0}\}$ and $V_{2x+1}=\{h(i):i\leq j_{1}\}$. By construction, $j_{0}\leq j_{1}\leq j_{0}+1$, and we have that $x\in\operatorname{ran}(f)$ if and only if $j_{1}=j_{0}+1$. ∎
@@
We prove (1); the proof of (2) is analogous. Fix $\langle X,\mathcal{U},k\rangle$ satisfying (1a)–(1d) in Definition 3.12. (In fact, we will not need to use (1d).) We define sequences $x_{0},x_{1},\ldots\in X$ and $U_{n_{0}},U_{n_{1}},\ldots\in\mathcal{U}$ such that $x_{i}\in U_{n_{j}}$ if and only if $i\leq j$. We claim that $Y=\{x_{0},x_{1},\ldots\}$ is a subspace of $X$ with the initial segment topology, as witnessed by the map $h:i\mapsto x_{i}$. Note that $U_{n_{j}}\cap Y=\{x_{i}:i\leq j\}$ for each $j$, so we wish to show that the basic open sets in $Y$ in addition to $\emptyset$ and $Y$ are precisely the sets $U_{n_{j}}\cap Y$ for some $j$. Since $U_{n_{j}}\in\mathcal{U}$, it is certainly true that $U_{n_{j}}\cap Y$ is basic open in $Y$. Conversely, consider any basic open set in $Y$ different from $\emptyset$ and $Y$. Then this equals $U\cap Y$ for some $U\in\mathcal{U}$ different from $\emptyset$ and $X$. For any $j$ such that $x_{j}\notin U$ it follows by property (1b) that $U\subseteq U_{n_{j}}$ and hence also that $U\subseteq U_{n_{j}}\smallsetminus\{x_{j}\}$. In particular, $U\cap Y\subseteq(U_{n_{j}}\smallsetminus\{x_{j}\})\cap Y$. We must therefore have $x_{0}\in U$ since $U\cap Y\neq\emptyset$ by assumption. Now since $Y\neq X$, it follows by property (1a) that $U$ is finite, and we can consequently fix the largest $j$ such that $x_{i}\in U$ for all $i\leq j$. We conclude that $U_{n_{j}}\cap Y\subseteq U\cap Y$. And since $x_{j+1}\notin U$, the preceding argument implies that $U\cap Y\subseteq(U_{n_{j+1}}\smallsetminus\{x_{j+1}\})\cap Y=U_{n_{j}}\cap Y$. Thus $U\cap Y=U_{n_{j}}\cap Y$, desired.
It remains to show how to construct the sequences $x_{0},x_{1},\ldots$ and $U_{n_{0}},U_{n_{1}},\ldots$, which we do by primitive recursion. Let $x_{0}$ be any element of $X$, and $U_{n_{0}}$ any element of $\mathcal{U}$ containing $x_{0}$. Now fix $j$, and suppose we have defined $x_{i}$ and $U_{n_{i}}$ for all $i\leq j$. By property (1b), and the fact that $x_{j}\in U_{n_{j}}\smallsetminus U_{n_{i}}$ for all $i<j$ by inductive hypothesis, we must have $U_{n_{i}}\subseteq U_{n_{j}}$ for all $i\leq j$. Let $s=|U_{j}|$. Using property (1c) let $U\in\mathcal{U}$ be a finite set with at least $s+1$ many elements. Since $\mathsf{I}\Sigma^{0}_{1}$ proves that there is no injection from a larger finite set into a strictly smaller one, we cannot have $U\subseteq U_{n_{j}}$. Hence, $U_{n_{j}}\subseteq U$ by property (1b), and this containment must be strict. Fixing any $x\in U\smallsetminus U_{n_{j}}$ thus means that $x_{i}\in U$ and $x\notin U_{i}$ for all $i\leq j$. Now by property (1a), being a finite element of $\mathcal{U}$ is definable by a $\Sigma^{0}_{1}$ formula (with parameters). Thus, the least $U$ and $x\in U$ satisfying this conclusion are definably by a $\Sigma^{0}_{1}$ formula, and we let $U_{n_{j+1}}=U$ and $x_{j+1}=x$. This completes the recursive definition and completes the proof. ∎
@@
Write $\mathcal{U}=\langle U_{n}:n\in\mathbb{N}\rangle$. Suppose $X$ has the initial segment topology; the case where it has the final segment topology is symmetric. Let $h:\mathbb{N}\to X$ be the witnessing homeomorphism, and consider any open code $O$. If $O_{\mathcal{U}}=X$ then it clearly belongs to $\mathcal{U}$, so suppose otherwise. Then every $U_{n}$ for $n\in\operatorname{ran}(O)$ is either $\emptyset$ or has the form $\{h(i):i\leq j_{n}\}$ for some $j_{n}$, and the numbers $j_{n}$ must be bounded else $O_{\mathcal{U}}$ would again be $X$. In $\mathsf{RCA}_{0}$, we can thus fix the least $m$ such that $j_{n}\leq j_{m}$ for all $n\in\operatorname{ran}(O)$ for which $j_{n}$ is defined. Then $O_{\mathcal{U}}=U_{m}\in\mathcal{U}$. ∎
@@ 4. General spaces
For the implication from (1) to (2), we note that the proof of the Ginsburg–Sands theorem given in Section 2 readily formalizes in $\mathsf{ACA}_{0}$ when applied to CSC spaces. The implication from (2) to (3) is trivial. It thus remains to show that (3) implies (1). To this end, we argue in $\mathsf{RCA}_{0}$ and assume $\mathsf{wGS}$. Note that $\mathsf{wGS}$ trivially implies $\mathsf{wGS}^{\operatorname{cl}}$, which we in turn prove is equivalent to $\mathsf{GST}_{1}$ in Theorem 4.8 below. Moreover, in Corollary 6.10 we show that $\mathsf{GST}_{1}$ implies $\mathsf{RT}^{2}_{2}$, and hence $\mathsf{B}\Sigma^{0}_{2}$. Therefore, we may avail ourselves of $\mathsf{B}\Sigma^{0}_{2}$ in our argument.
We prove that the range of a given one-to-one function $f:\mathbb{N}\to\mathbb{N}$ exists. Given $x,s\in\mathbb{N}$, using the notation from the proof of Proposition 3.10, define
@@
and let $\langle\mathbb{N},\mathcal{U},k\rangle$ be the CSC space generated by $\langle V_{n}:n\in\mathbb{N}\rangle$.
@@
We claim that if a number $x$ is sufficiently large with respect to a number $w$ then $w$ belongs to every basic open set containing $x$. To see this, fix $w$ and choose $x>w$ so that $\operatorname{ran}(f)\operatorname{\upharpoonright}w=\operatorname{ran}(f)\operatorname{\upharpoonright}w~{}[x]$, which exists by $\mathsf{B}\Sigma^{0}_{2}$. Fix $U\in\mathcal{U}$ containing $x$. Then $U$ is a finite intersection of sets of the form $V_{\langle z,s\rangle}$ for some $z$ and $s$, and since no element of $V_{\langle z,s\rangle}$ can be larger than $z$, we necessarily have $z\geq x$. Consider such a $V_{\langle z,s\rangle}$. If $x=z$, then $w\in V_{\langle z,s\rangle}=V_{\langle x,s\rangle}$ since $\operatorname{ran}(f)\operatorname{\upharpoonright}w~{}[t]=\operatorname{ran}(f)\operatorname{\upharpoonright}w$ for all $t\geq x$ by choice of $x$. If $x<z$ then we must have $\operatorname{ran}(f)\operatorname{\upharpoonright}x~{}[z]=\operatorname{ran}(f)\operatorname{\upharpoonright}x~{}[z+s]$, and hence also $\operatorname{ran}(f)\operatorname{\upharpoonright}w~{}[z]=\operatorname{ran}(f)\operatorname{\upharpoonright}w~{}[z+s]$ since $w<x$, so $w\in V_{\langle z,s\rangle}$. This proves the claim.
Apply $\mathsf{wGS}$ to $X$ to obtain an infinite subspace $Y$. By the preceding claim, we immediately know that $Y$ cannot be $T_{1}$ nor have the final segment topology. And since, for $w<x$, we have that $V_{\langle w,0\rangle}$ contains $w$ but not $x$, we know that $Y$ cannot be indiscrete. This means that $Y$ has the initial segment topology. There is thus a bijection $h:\mathbb{N}\to Y$ such that the basic open sets of $Y$ in the subspace topology are precisely $\{h(i):i\leq j\}$ for some $j\in\mathbb{N}$. Let $Y=\{y_{0}<y_{1}<\cdots\}$; we claim that $h(n)=y_{n}$ for all $n$. To see this, first suppose that $h(n)>h(m)$ for some $n<m$, and fix $j$ such that $V_{\langle h(m),0\rangle}\cap Y=\{h(i):i\leq j\}$. Then $V_{\langle h(m),0\rangle}\cap Y$ contains $h(n)$, contradicting the fact that this intersection only contains $h(m)$ and otherwise numbers smaller than $y_{m}$. It follows that $h(n)<h(m)$ for all $n<m$, whence an easy induction using the fact that $h$ is bijective shows that $h(n)=y_{n}$. This proves the claim. In particular, every basic open set in $X$ is closed downward in $Y$ under $\leq$.
We now use $Y$ to prove the existence of $\operatorname{ran}(f)$. Fix $w\in\mathbb{N}$. Since $Y$ is infinite, we can find $z>x>w$ with $x,z\in Y$. We must then have that $\operatorname{ran}(f)\operatorname{\upharpoonright}x=\operatorname{ran}(f)\operatorname{\upharpoonright}x~{}[z]$, and hence also that $\operatorname{ran}(f)\operatorname{\upharpoonright}w=\operatorname{ran}(f)\operatorname{\upharpoonright}w~{}[z]$. If not, then there would be an $s$ such that $\operatorname{ran}(f)\operatorname{\upharpoonright}x~{}[z]\neq\operatorname{ran}(f)\operatorname{\upharpoonright}x~{}[z+s]$. But then $V_{\langle z,s\rangle}$ would be a basic open set in $X$ containing $z$ but not $x<z$, even though both belong to $Y$. This completes the proof. ∎